OpenStudy (anonymous):

Heyy guys can you help me pleasee???? Can you tell me how to find the quadratic equation for (4,-1) ?? That is the vertex point. I need to find the roots, Please help!! :[

OpenStudy (anonymous):

you need more than just one point to find the roots if you just have the vertex that is not enough

OpenStudy (anonymous):

do you have any other piece of information? two more points? a picture?

OpenStudy (anonymous):

i really dont remember /: uh dont you have to use the equation (x-h)^2+k??

OpenStudy (raden):

y = a(x-h)^2+k with a=constant, h,k is vertex so, to find a, u need an extra point coordinat again

OpenStudy (anonymous):

i think i got it now guys, thanks thou, i really appreciate it

OpenStudy (anonymous):

you can use \(y=a(x-4)^2-1\) but this will not help you find the zeros because you do not know what \(a\) is without some other point known to you